# AI-Face: A Million-Scale Demographically Annotated AI-Generated Face Dataset and Fairness Benchmark

### v1

@@ B.1 Details of Detection Methods
Table 15: Summary of the implemented detectors in our fairness benchmark.
@@
B.2 Implementation Details
For fairness benchmark, all experiments are based on the PyTorch with a single NVIDIA RTX A6000 GPU. During training, we utilize SGD optimizer with a learning rate of 0.0005, with momentum of 0.9 and weight decay of 0.005. The batch size is set to 128 for most detectors. However, for the SRM [66], UCF [16], and PG-FDD [21], the batch size is adjusted to 32 due to GPU memory. For hyperparameters defined in these detectors, we use the default values set in their original papers. All detectors are initialized with their official pre-trained weights, and trained for 5 epochs.
B.3 Fairness Metrics
We assume a test set comprising indices {1, …, $n$}. $Y_{j}$ and $\hat{Y}_{j}$ respectively represent the true and predicted labels of the sample $X_{j}$. Their values are binary, where 0 means real and 1 means fake. For all fairness metrics, a lower value means better performance.
@@ B.3 Fairness Metrics
$F_{IND}:=\sum_{j=1}^{n}\sum_{l=j+1}^{n}\mathbb{I}_{[\left|f(X_{j})-f(X_{l})\right|-\delta\|X_{j}-X_{l}\|]},$
[…]
B.4 Full Subsets Evaluation Results
@@
Detailed test results of each subset as shown from Table 16 to Table 35 are presented in this section. The findings align with the results reported in Fig. 4.
Model Type
Naive Frequency Spatial Fairness-enhanced
Dataset Attribute Metric
@@
FF++ Gender $F_{MEO}$ 4.353 3.346 1.161 2.595 0.887 2.492 4.916 10.873 2.516 12.198 1.606 2.214
$F_{DP}$ 1.250 1.096 0.276 0.601 0.392 0.409 1.231 2.874 0.61 2.722 1.024 0.772
$F_{OAE}$ 0.177 0.132 0.396 0.426 0.231 0.228 0.489 0.015 0.196 0.977 0.941 0.095
$F_{EO}$ 4.749 4.293 1.335 2.728 1.012 2.839 5.117 12.323 3.221 12.993 2.969 2.362
Race $F_{MEO}$ 10.304 9.813 7.630 15.051 6.844 22.26 9.791 23.588 4.564 22.598 2.607 15.657
$F_{DP}$ 3.562 9.544 3.485 3.22 5.864 3.516 5.554 12.934 8.75 8.954 6.65 2.943
$F_{OAE}$ 4.465 7.396 6.232 4.045 2.541 5.227 4.388 6.889 3.522 7.382 1.939 3.764
$F_{EO}$ 17.066 27.404 12.835 20.586 11.277 36.221 17.288 70.499 11.944 48.644 6.882 18.386
Age $F_{MEO}$ 9.851 5.348 5.984 6.204 3.622 14.005 9.692 15.205 9.423 24.413 1.857 6.136
$F_{DP}$ 2.887 4.708 1.280 5.661 6.479 7.919 6.196 9.205 7.693 4.346 6.221 4.339
$F_{OAE}$ 1.038 5.813 6.417 2.049 0.856 1.581 2.606 8.927 1.138 4.263 1.472 2.112
$F_{EO}$ 18.191 11.876 8.199 13.665 5.636 20.696 17.291 29.607 14.781 47.613 6.419 12.446
Intersection $F_{MEO}$ 28.949 16.662 11.994 18.672 8.505 30.828 19.132 54.201 8.784 39.858 5.130 16.994
$F_{DP}$ 11.648 12.215 5.127 6.721 10.157 4.449 11.268 32.584 14.697 20.864 10.087 4.831
$F_{OAE}$ 8.442 10.876 10.295 7.210 4.868 8.742 5.638 15.415 8.209 10.843 3.322 4.491
$F_{EO}$ 70.162 68.005 32.625 48.53 25.922 78.296 40.971 169.535 33.428 131.755 19.399 38.887
- ACC 92.280 89.282 86.051 94.832 93.676 92.587 94.982 83.652 95.183 91.420 93.254 96.237
AUC 95.605 91.281 83.542 97.878 97.820 96.164 98.115 76.839 98.147 94.618 97.996 98.245
AP 99.207 98.381 96.712 99.631 99.619 99.29 99.668 95.321 99.684 99.011 99.658 99.681
EER 10.951 16.807 24.299 6.756 6.565 9.888 6.02 30.755 5.993 12.449 6.429 7.273
Table 16: Detailed fairness and utility evaluation results on FF++.
Model Type
Naive Frequency Spatial Fairness-enhanced
Dataset Attribute Metric
@@
DFDC Gender $F_{MEO}$ 6.011 3.708 5.567 4.415 2.357 8.711 1.492 7.444 2.062 4.87 2.944 3.687
$F_{DP}$ 5.878 0.998 3.43 4.959 3.829 8.35 3.776 5.039 3.662 5.271 4.468 6.348
$F_{OAE}$ 2.97 2.744 1.88 2.645 2.537 2.438 1.427 0.075 2.178 3.012 2.233 0.869
$F_{EO}$ 6.222 5.564 7.742 5.609 3.833 10.841 2.517 12.51 3.784 4.898 3.95 4.968
Race $F_{MEO}$ 8.525 6.846 24.319 7.667 9.726 9.139 11.603 22.342 10.74 15.529 11.403 5.992
$F_{DP}$ 21.619 11.534 20.596 25.03 24.594 21.463 25.9 24.317 26.634 23.997 25.534 24.613
$F_{OAE}$ 1.622 3.701 12.756 3.048 2.816 5.793 4.722 11.051 2.699 14.659 3.46 2.523
$F_{EO}$ 26.728 15.611 47.784 25.2 25.744 24.09 22.014 65.788 23.784 47.679 26.64 12.268
Age $F_{MEO}$ 6.193 7.721 17.868 5.022 7.375 10.382 4.608 13.078 5.683 20.119 5.578 3.96
$F_{DP}$ 11.068 4.752 14.277 9.967 12.117 8.172 11.987 9.764 9.112 13.229 10.702 11.48
$F_{OAE}$ 2.817 5.951 4.984 3.918 2.585 6.092 2.513 7.523 3.869 12.581 2.498 1.653
$F_{EO}$ 14.397 16.233 26.396 14.327 13.88 16.625 12.03 22.816 11.274 31.018 8.736 6.954
Intersection $F_{MEO}$ 14.479 15.029 33.979 14.067 24.924 14.117 16.119 38.533 17.421 20.447 18.268 10.973
$F_{DP}$ 28.877 17.816 30.153 32.117 31.493 27.666 31.604 28.815 33.791 27.812 30.224 31.389
$F_{OAE}$ 5.619 8.088 20.771 4.456 7.453 9.306 5.922 14.994 4.423 18.877 5.642 3.832
$F_{EO}$ 72.695 60.893 111.03 59.238 67.19 60.749 63.262 133.283 58.174 90.761 64.155 33.495
- ACC 81.223 71.939 71.044 87.658 87.482 83.536 89.155 64.164 88.75 81.452 88.867 92.905
AUC 90.395 80.17 81.942 95.158 95.789 91.837 96.025 72.228 95.65 91.695 95.916 97.014
AP 91.284 81.442 82.547 95.764 96.313 92.435 96.567 75.304 96.219 92.37 96.447 97.081
EER 18.443 28.133 26.271 12.367 10.805 15.588 10.818 33.542 10.927 17.043 10.709 8.317
Table 17: Detailed fairness and utility evaluation results on DFDC.
Model Type
Naive Frequency Spatial Fairness-enhanced
Dataset Attribute Metric
@@
DFD Gender $F_{MEO}$ 7.052 2.139 3.857 6.07 0.269 10.037 6.095 1.039 2.257 3.605 3.059 0.717
$F_{DP}$ 5.543 5.864 1.231 4.871 7.154 2.188 5.342 2.327 7.261 6.893 8.593 6.827
$F_{OAE}$ 3.445 3.199 5.624 2.212 0.232 4.177 2.313 5.785 2.198 2.996 2.609 1.241
$F_{EO}$ 8.657 3.812 3.868 6.325 0.467 10.381 6.731 1.300 4.139 5.855 3.900 1.326
Race $F_{MEO}$ 5.975 6.844 12.306 5.574 0.319 18.91 6.141 20.641 6.292 10.597 6.021 11.641
$F_{DP}$ 25.863 19.116 11.976 25.678 40.64 21.081 28.174 14.104 26.949 28.784 28.439 29.743
$F_{OAE}$ 6.002 10.714 16.754 4.602 0.206 8.565 3.89 15.842 4.917 4.594 4.125 3.797
$F_{EO}$ 16.002 17.914 24.819 14.628 0.884 32.788 15.477 51.098 17.872 19.959 13.855 17.467
Age $F_{MEO}$ 14.485 13.629 2.744 9.24 0.9 9.38 10.383 6.69 10.768 10.107 10.942 5.6
$F_{DP}$ 34.386 18.578 10.063 32.355 20.119 18.826 33.553 4.892 34.865 32.253 34.503 27.165
$F_{OAE}$ 11.001 18.255 22.847 6.943 0.434 13.797 7.41 23.7 5.315 7.635 6.256 6.095
$F_{EO}$ 22.487 33.616 6.473 15.786 1.97 13.896 18.326 12.859 16.44 14.035 13.272 8.349
Intersection $F_{MEO}$ 15.691 37.9 20.833 13.62 1.786 27.246 11.053 35.828 18.056 20.833 9.157 12.903
$F_{DP}$ 35.824 31.581 18.295 36.56 53.771 29.097 38.828 28.054 38.536 41.172 39.027 41.388
$F_{OAE}$ 9.913 15.939 21.972 6.863 1.322 11.216 6.327 22.706 7.158 6.101 6.097 5.31
$F_{EO}$ 46.408 79.825 68.93 42.743 7.073 91.155 41.325 111.273 53.592 49.779 40.678 41.822
- ACC 93.039 88.321 83.862 94.6 99.505 91.405 94.984 80.753 94.761 92.99 94.6 97.102
AUC 97.507 93.914 89.886 98.478 99.942 96.347 98.592 82.817 98.651 97.659 98.813 99.082
AP 99.349 98.366 97.059 99.596 99.965 98.929 99.614 95.008 99.62 99.375 99.687 99.75
EER 8.086 13.377 18.014 6.183 0.500 10.048 6.124 24.911 5.945 7.788 5.529 5.470
Table 18: Detailed fairness and utility evaluation results on DFD.
Model Type
Naive Frequency Spatial Fairness-enhanced
Dataset Attribute Metric

### v3

@@ C.2 Details of Detection Methods
Table C.1: Summary of the implemented detectors in our fairness benchmark.
@@
C.3 Fairness Metrics
We assume a test set comprising indices {1, …, $n$}. $Y_{j}$ and $\hat{Y}_{j}$ respectively represent the true and predicted labels of the sample $X_{j}$. Their values are binary, where 0 means real and 1 means fake. For all fairness metrics, a lower value means better performance. The formulations of fairness metrics are as follows,
@@ C.3 Fairness Metrics
$F_{IND}:=\sum_{j=1}^{n-1}\sum_{l=j+1}^{n}[\left|f(X_{j})-f(X_{l})\right|-\delta\|X_{j}-X_{l}\|_{2}]_{+},$
where $D$ is the demographic variable, $\mathcal{J}$ is the set of subgroups with each subgroup $\mathcal{J}_{j}\in\mathcal{J}$. $M$ is the set of detection models and $F$ is the set of fairness metrics. $F_{EO}$ measures the disparity in TPR and FPR between each subgroup and the overall population. $F_{OAE}$ measures the maximum ACC gap across all demographic groups. $F_{DP}$ measures the maximum difference in prediction rates across all demographic groups. And $F_{MEO}$ captures the largest disparity in prediction outcomes (either positive or negative) when comparing different demographic groups. $\delta$ in $F_{IND}$ is a predefined scale factor (0.08 in our experiments). $[\cdot]_{+}$ is the hinge function, $\|\cdot\|_{2}$ is the $\ell_{2}$ norm. $f(X_{j})$ represents the predicted logits of the model for input sample $X_{j}$. $F_{IND}$ points that a model should be fair across individuals if similar individuals have similar predicted outcomes.
Appendix D More Fairness Benchmark Results and Analysis
D.1 Detailed Results of Overall Performance Comparison
@@
Detailed test results of each subgroup of each detector on AI-Face are presented in this section. Table D.1 provides comprehensive metrics of each subgroup on AI-Face. These results and findings align with the results reported in Table. 4 submitted main manuscript.
Model Type Method Metric Gender Skin Tone Age Intersection
M F L M D Child Young Adult Middle Senior M-L M-M M-D F-L F-M F-D
Naive Xception  [82] AUC 98.90 98.20 97.69 98.44 98.88 95.95 97.86 99.10 98.66 96.54 97.88 98.70 99.22 97.53 98.17 98.19
FPR 11.12 15.10 18.37 14.72 9.54 36.83 18.26 8.95 12.78 20.42 17.70 12.82 8.42 18.93 16.58 11.32
TPR 99.38 99.22 98.94 99.62 98.43 99.55 99.38 99.21 99.49 98.51 98.55 99.64 98.87 99.16 99.59 97.58
ACC 96.77 95.80 95.16 96.43 96.03 89.83 94.98 97.10 97.06 92.06 94.26 96.79 96.78 95.70 96.10 94.67
EfficientB4  [83] AUC 98.86 98.31 99.23 98.94 97.59 99.63 98.61 98.44 98.82 98.69 99.01 99.10 98.36 99.39 98.78 95.94
FPR 17.57 22.96 21.32 17.17 25.47 16.13 22.92 19.79 19.41 19.07 20.50 15.44 20.74 22.00 18.87 33.02
TPR 99.04 98.56 99.49 99.02 98.19 99.56 98.89 98.55 99.14 98.66 99.22 99.25 98.60 99.64 98.81 97.40
ACC 94.91 93.42 94.95 95.43 91.05 95.37 93.45 93.81 95.48 92.62 94.01 95.90 93.04 95.51 94.99 87.44
ViT-B/16  [84] AUC 99.02 98.26 97.50 98.77 98.49 95.55 98.23 98.98 98.80 97.28 97.21 99.03 99.03 97.80 98.50 97.21
FPR 14.06 19.17 21.74 16.86 15.48 28.28 21.09 13.34 17.22 20.92 21.43 14.93 12.61 22.00 18.76 20.05
TPR 98.43 97.51 96.75 98.23 97.36 94.58 97.88 98.14 98.40 96.16 96.77 98.56 98.21 96.74 97.93 95.72
ACC 95.33 93.52 92.72 94.88 93.49 88.48 93.16 95.17 95.31 90.34 91.96 95.48 95.10 93.16 94.33 90.55
Frequency F3Net  [85] AUC 99.09 98.24 98.51 98.68 98.79 96.05 98.17 99.15 98.97 97.40 98.66 99.00 99.27 98.30 98.36 97.80
FPR 12.49 17.21 30.72 16.54 10.78 49.32 21.21 10.51 12.52 20.00 30.75 14.33 9.38 30.69 18.72 13.02
TPR 99.15 99.00 99.69 99.41 98.11 99.77 99.40 99.00 99.06 97.96 99.55 99.34 98.76 99.76 99.48 96.86
ACC 96.25 95.12 93.05 95.87 95.43 86.66 94.27 96.54 96.77 91.84 91.55 96.22 96.42 93.95 95.55 93.63
SPSL  [86] AUC 98.88 98.58 98.99 98.70 98.81 97.41 98.05 99.17 98.75 97.32 99.05 98.78 99.02 98.92 98.61 98.36
FPR 11.62 16.03 19.50 14.77 11.44 37.20 20.20 9.58 13.12 19.07 18.94 12.88 9.57 19.95 16.62 14.43
TPR 99.64 99.52 99.53 99.73 99.15 99.78 99.58 99.53 99.69 99.12 99.44 99.77 99.37 99.58 99.70 98.73
ACC 96.84 95.80 95.38 96.51 95.96 89.90 94.65 97.17 97.16 92.92 94.59 96.88 96.80 95.85 96.17 94.42
SRM  [87] AUC 98.45 97.40 98.71 97.94 97.95 97.24 97.27 98.42 98.13 97.78 99.24 98.46 98.52 98.36 97.48 96.82
FPR 12.84 19.11 22.30 17.50 12.30 36.76 22.42 11.92 15.33 19.64 18.94 14.70 9.92 25.06 20.25 16.10
TPR 98.84 98.33 99.41 99.02 97.35 99.23 98.91 98.29 98.97 97.52 99.89 99.20 98.06 99.16 98.85 95.99
ACC 95.93 94.16 94.67 95.35 94.44 89.62 93.59 95.65 96.14 91.67 94.91 96.03 95.76 94.53 94.72 92.03
Spatial UCF  [26] AUC 98.62 97.45 97.20 97.92 98.49 95.59 97.26 98.74 98.67 97.04 97.67 98.44 99.00 96.88 97.41 97.47
FPR 11.30 16.37 26.23 16.01 8.90 55.93 21.10 8.43 11.70 20.40 24.85 13.82 7.23 27.37 18.16 11.57
TPR 98.20 97.77 98.75 98.37 96.89 99.53 98.50 97.58 98.35 96.91 99.00 98.47 97.63 98.61 98.28 95.45
ACC 95.84 94.39 93.30 95.18 95.14 84.71 93.61 96.03 96.36 91.01 92.70 95.66 96.24 93.65 94.73 93.15
UnivFD  [88] AUC 98.55 97.76 98.38 98.47 97.40 98.13 98.07 98.33 98.11 96.94 98.24 98.76 98.13 98.46 98.19 95.84
FPR 16.46 20.97 20.06 19.04 17.60 19.90 19.91 16.70 22.13 17.04 17.70 16.84 15.88 22.00 21.20 20.36
TPR 98.02 97.12 97.57 98.26 95.69 97.43 97.40 97.39 98.41 94.49 97.10 98.54 96.99 97.83 98.02 93.17
ACC 94.42 92.80 93.73 94.43 91.68 92.80 93.09 93.74 94.35 90.56 93.19 95.03 93.29 94.04 93.87 88.74
CORE  [89] AUC 99.04 98.01 97.80 98.47 98.79 95.88 97.82 99.09 98.77 97.11 98.12 98.91 99.26 97.57 98.02 97.84
FPR 10.73 16.52 21.46 14.76 10.68 43.07 19.42 9.19 12.34 20.10 18.63 12.14 8.45 23.79 17.34 14.25
TPR 99.40 99.27 99.61 99.51 98.84 99.83 99.27 99.27 99.46 99.00 99.78 99.53 99.13 99.52 99.49 98.28
ACC 96.88 95.49 95.01 96.34 95.97 88.37 94.61 97.08 97.13 92.49 94.91 96.87 96.95 95.07 95.85 94.17
Fairness- enhanced DAW-FDD  [29] AUC 98.36 97.15 96.84 97.45 98.51 94.97 96.41 98.62 98.00 94.95 96.83 98.05 98.85 96.77 96.87 97.91
FPR 14.12 19.63 26.93 18.59 12.81 56.69 23.64 11.23 15.40 26.17 26.40 16.16 10.80 27.37 20.98 16.03
TPR 99.42 99.24 99.73 99.38 99.20 99.77 99.18 99.28 99.51 98.98 99.78 99.47 99.32 99.70 99.29 98.97
ACC 96.05 94.73 93.91 95.39 95.58 84.69 93.49 96.56 96.56 90.41 92.86 95.90 96.41 94.53 94.91 94.06
DAG-FDD  [29] AUC 99.05 98.44 97.56 98.79 98.73 96.55 98.14 99.10 98.91 97.59 98.11 99.00 99.16 97.10 98.57 97.83
FPR 12.11 18.02 20.76 15.10 14.21 26.95 20.01 11.72 14.61 22.40 18.32 13.04 10.58 22.76 17.14 20.00
TPR 99.21 99.05 98.98 99.25 98.83 97.64 98.92 99.18 99.37 99.13 98.77 99.33 98.98 99.10 99.17 98.54
ACC 96.39 94.97 94.67 96.06 94.90 91.07 94.20 96.36 96.61 91.79 94.26 96.51 96.23 94.92 95.65 92.47
PG-FDD  [30] AUC 99.36 98.94 98.94 99.18 99.13 98.83 98.89 99.35 99.27 97.85 98.97 99.35 99.39 98.89 99.02 98.51
FPR 9.49 12.68 15.29 12.06 8.82 22.77 14.31 7.97 11.64 19.74 13.35 10.90 7.30 16.88 13.20 11.26
TPR 98.73 98.21 98.63 98.85 97.43 99.11 98.44 98.28 98.87 97.74 98.66 99.02 98.12 98.61 98.69 96.10
ACC 96.68 95.61 95.59 96.43 95.55 93.27 95.26 96.66 96.79 91.78 95.49 96.75 96.56 95.65 96.12 93.69
Table D.1: Detailed test results of each subgroup of each detector on the AI-Face. In the Skin Tone groups, ‘L’ represents Light (Tone 1-3), ‘M’ is Medium (4-6), ‘D’ is Dark (Tone 7-10).
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced
@@
Fairness(%) Skin Tone $F_{MEO}$ 10.901 4.384 17.219 14.583 9.620 15.508 14.978 2.441 13.135 12.519 12.597 13.965
$F_{DP}$ 11.274 9.191 10.713 12.739 11.711 11.282 11.549 8.117 11.968 11.179 12.026 11.768
$F_{OAE}$ 2.434 3.609 2.276 2.232 2.814 1.780 1.950 2.940 1.658 0.878 1.753 1.439
$F_{EO}$ 0.160 0.093 0.205 0.209 0.156 0.191 0.186 0.034 0.176 0.165 0.176 0.176
Gender $F_{MEO}$ 5.475 5.458 8.003 5.749 5.754 5.848 5.575 3.244 4.367 5.186 5.808 4.086
$F_{DP}$ 1.205 1.412 2.416 1.340 1.445 1.959 1.810 0.781 0.980 1.715 1.470 1.545
$F_{OAE}$ 2.043 1.800 1.896 2.054 1.969 1.471 1.569 1.413 1.848 1.430 2.012 1.133
$F_{EO}$ 0.066 0.063 0.083 0.068 0.067 0.062 0.060 0.043 0.056 0.056 0.068 0.044
Age $F_{MEO}$ 28.244 7.460 38.521 27.860 24.768 40.542 44.342 8.584 34.156 36.450 35.031 36.197
$F_{DP}$ 11.228 12.245 12.140 11.395 11.466 14.564 15.856 16.134 13.525 12.256 13.478 12.082
$F_{OAE}$ 7.138 5.234 10.940 6.933 6.053 11.126 11.481 4.171 8.294 9.192 8.636 8.934
$F_{EO}$ 0.460 0.175 0.560 0.460 0.410 0.550 0.610 0.191 0.508 0.537 0.524 0.539
Intersection $F_{MEO}$ 15.752 10.644 24.460 18.455 15.157 18.381 17.397 5.300 16.257 14.806 15.219 16.517
$F_{DP}$ 16.943 14.565 13.773 18.071 17.490 14.943 15.612 12.967 17.063 14.802 16.786 15.513
$F_{OAE}$ 6.805 8.029 5.025 6.658 7.200 3.614 3.532 6.226 5.079 3.314 5.757 2.989
$F_{EO}$ 0.355 0.307 0.441 0.440 0.366 0.382 0.382 0.178 0.371 0.336 0.399 0.354
Utility(%) - AUC 0.968 0.968 0.981 0.966 0.968 0.967 0.977 0.979 0.975 0.970 0.973 0.978
ACC 0.931 0.922 0.924 0.930 0.929 0.946 0.951 0.933 0.947 0.941 0.941 0.952
AP 0.987 0.988 0.994 0.986 0.988 0.985 0.991 0.993 0.989 0.987 0.989 0.991
EER 0.093 0.101 0.082 0.096 0.095 0.076 0.074 0.083 0.076 0.085 0.084 0.072
FPR 0.205 0.219 0.290 0.199 0.190 0.205 0.163 0.188 0.144 0.186 0.168 0.151
Table D.2: Detailed fairness and utility evaluation results on 20% training subset.
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced
@@
Fairness(%) Skin Tone $F_{MEO}$ 9.815 4.414 12.194 10.801 9.275 24.037 12.299 2.568 15.734 12.628 10.463 10.982
$F_{DP}$ 10.080 10.413 9.475 11.137 10.644 12.550 10.632 8.322 11.251 10.157 10.456 10.928
$F_{OAE}$ 0.122 0.095 0.146 0.154 0.122 0.281 0.150 0.043 0.180 0.154 0.135 0.143
$F_{EO}$ 1.472 3.796 3.395 2.088 1.631 3.280 1.533 2.898 2.384 1.571 1.323 1.188
Gender $F_{DP}$ 5.576 3.592 6.089 5.985 4.368 7.959 4.400 3.438 5.234 5.960 4.797 5.390
$F_{MEO}$ 1.817 0.822 1.853 1.566 1.227 2.658 1.481 0.797 1.866 2.053 1.458 2.052
$F_{OAE}$ 1.559 1.622 1.829 1.966 1.595 1.722 1.338 1.507 1.369 1.473 1.530 1.303
$F_{EO}$ 0.060 0.047 0.067 0.069 0.052 0.080 0.049 0.045 0.055 0.062 0.055 0.055
Age $F_{MEO}$ 32.781 9.931 18.050 26.665 33.004 54.967 43.829 7.840 38.202 41.707 34.285 33.582
$F_{DP}$ 12.161 12.428 14.762 10.954 13.272 16.006 14.243 16.076 13.394 13.661 11.630 12.955
$F_{OAE}$ 8.535 4.312 6.949 6.831 8.210 14.795 11.248 4.185 9.906 10.841 8.738 8.520
$F_{EO}$ 0.474 0.235 0.348 0.432 0.470 0.738 0.611 0.179 0.541 0.585 0.509 0.500
Intersection $F_{MEO}$ 13.451 8.558 15.615 13.559 12.356 30.133 15.278 5.585 19.342 17.006 12.784 13.197
$F_{DP}$ 13.795 16.478 14.409 16.096 15.138 15.893 14.797 13.263 15.123 14.586 14.345 15.166
$F_{OAE}$ 4.424 7.462 5.221 6.173 4.728 5.030 2.886 6.298 3.741 3.541 4.402 2.911
$F_{EO}$ 0.312 0.236 0.416 0.360 0.309 0.550 0.304 0.185 0.374 0.324 0.310 0.301
Utility(%) - AUC 0.979 0.976 0.980 0.974 0.982 0.957 0.979 0.979 0.981 0.976 0.978 0.982
ACC 0.951 0.940 0.933 0.937 0.951 0.938 0.955 0.934 0.957 0.948 0.949 0.960
AP 0.991 0.991 0.993 0.989 0.993 0.983 0.992 0.993 0.991 0.989 0.991 0.992
EER 0.068 0.078 0.081 0.083 0.066 0.117 0.072 0.082 0.058 0.074 0.074 0.055
FPR 0.165 0.147 0.136 0.186 0.147 0.245 0.157 0.180 0.151 0.184 0.169 0.130
Table D.3: Detailed fairness and utility evaluation results on 40% training subset.
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced
@@
Fairness(%) Skin Tone $F_{MEO}$ 9.086 14.704 4.388 15.303 6.813 14.516 14.952 2.186 9.689 13.488 9.672 4.108
$F_{DP}$ 10.232 11.784 7.714 11.225 9.979 14.909 13.116 8.004 10.666 10.844 10.054 8.575
$F_{OAE}$ 1.531 2.017 2.572 2.320 1.247 1.733 1.562 2.777 1.208 1.664 1.259 1.383
$F_{EO}$ 0.124 0.194 0.084 0.177 0.100 0.234 0.208 0.043 0.131 0.163 0.125 0.055
Gender $F_{MEO}$ 4.418 6.743 8.445 5.545 5.242 7.331 5.713 4.182 4.395 5.579 4.978 2.430
$F_{DP}$ 1.243 2.063 2.697 2.096 1.736 3.656 1.845 1.142 1.622 2.153 1.600 1.061
$F_{OAE}$ 1.567 1.846 2.052 1.318 1.499 0.630 1.651 1.489 1.217 1.267 1.470 0.858
$F_{EO}$ 0.053 0.072 0.086 0.057 0.057 0.087 0.062 0.050 0.048 0.056 0.055 0.029
Age $F_{MEO}$ 35.231 27.998 17.573 42.366 35.428 37.043 34.243 5.520 28.666 40.326 38.409 24.355
$F_{DP}$ 12.874 12.663 11.691 15.070 12.924 16.570 13.579 15.256 11.264 13.929 13.112 10.449
$F_{OAE}$ 8.954 7.379 7.004 10.921 8.915 8.078 8.232 3.900 7.519 10.379 9.909 6.870
$F_{EO}$ 0.514 0.411 0.320 0.585 0.519 0.528 0.520 0.134 0.438 0.570 0.558 0.362
Intersection $F_{MEO}$ 11.554 18.923 10.063 18.907 10.175 20.404 18.818 5.414 12.995 15.944 12.552 5.425
$F_{DP}$ 14.625 15.884 9.908 15.240 14.093 18.967 18.087 12.890 15.331 14.584 13.949 13.379
$F_{OAE}$ 4.755 4.997 5.459 3.617 4.210 2.965 4.832 6.152 3.747 3.106 4.102 3.306
$F_{EO}$ 0.279 0.405 0.311 0.362 0.251 0.477 0.431 0.200 0.276 0.331 0.282 0.159
Utility(%) - AUC 0.976 0.980 0.986 0.981 0.983 0.976 0.981 0.981 0.982 0.977 0.982 0.982
ACC 0.948 0.945 0.943 0.961 0.952 0.927 0.951 0.935 0.956 0.960 0.950 0.960
AP 0.989 0.992 0.996 0.991 0.993 0.991 0.992 0.994 0.992 0.987 0.993 0.991
EER 0.074 0.071 0.065 0.058 0.067 0.092 0.068 0.078 0.062 0.060 0.069 0.058
FPR 0.166 0.183 0.177 0.137 0.154 0.137 0.137 0.198 0.145 0.143 0.172 0.127
Table D.4: Detailed fairness and utility evaluation results on 60% training subset.
Measure Attribute Metric Model Type
Native Frequency Spatial Fairness-enhanced